# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 30 Mar 2026 v3

@@ Abstract
The prim-lwe problem, introduced by Sehrawat, Yeo, and Desmedt (Theoretical Computer Science, 886 (2021)), is a variant of the Learning with Errors problem in which the secret matrix is required to have a primitive-root determinant. The dimension-uniform reduction constant is $c(p)=\inf_{n\geq 1}c_{n}(p)$, where $c_{n}(p)$ is the exact density of $n\times n$ matrices over $\mathbb{F}_{p}$ with primitive-root determinant. Sehrawat, Yeo, and Desmedt asked whether $\inf_{p\text{ prime}}c(p)=0$, observing that an affirmative answer would follow from the conjectural infinitude of primorial primes. We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, bypassing the primorial-prime hypothesis. We further establish the sharp order
$\min_{p\leq x}c(p)\asymp\frac{1}{\log\log x}\qquad(x\to\infty),$
and show that the limiting distribution of $c(p)$ over the primes has support exactly $[0,1/2]$. This full-support property appears to be new. The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$ via a transport lemma, is moreover continuous and purely singular, and the limiting measure has Hausdorff dimension 0. The moments of the limiting distribution are given by convergent Euler products that uniquely determine it, and the associated Mellin transform extends analytically to the right half-plane. We also derive explicit lower bounds on $c(q)$ for primes of cryptographic interest, parameterized solely by the number of distinct prime factors of q1q-1. As a simple conservative explicit bound, for any prime $q>2^{30}$ the dimension-uniform expected overhead $1/c(q)$ is at most $1.79\log q$. On the other hand, our results show that the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$, and in particular $1/c(q)=O(\log\log q)$ pointwise.
@@ 1 Introduction
Let $p$ be a prime and $M_{n}(\mathbb{F}_{p})$ denote the ring of $n\times n$ matrices over $\mathbb{F}_{p}$. Following [1], a matrix $M\in M_{n}(\mathbb{F}_{p})$ has primitive-root determinant if $\det(M)$ generates $\mathbb{F}_{p}^{*}$. The fraction of such matrices is denoted $c_{n}(p)$; since it is decreasing in $n$, the dimension-uniform constant governing the reduction from decision-LWE to decision-prim-lwe is $c(p):=\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)$. The precise formulas are recorded in Section 2. Sehrawat, Yeo, and Desmedt [1, Remark 3(ii)] asked whether $\inf_{p}c(p)=0$ can be proved unconditionally; see Section 2 for the precise statement.
@@
We resolve this question and establish several natural refinements.
@@
2. Sharp order of the minimum (Proposition 1, Corollary 8). By additionally invoking Linnik’s theorem on the least prime in an arithmetic progression, together with standard prime-number-theorem consequences for primorials, we establish
@@
The matching lower bound follows from two observations: if $p\leq x$ is an odd prime, then $\omega(p-1)\leq m(x):=\max\{k:N_{k}\leq x\}$, and Mertens’ theorem converts this into a $1/\!\log\log x$ lower bound for $\varphi(p-1)/(p-1)$; the additional factor $\prod_{j\geq 1}(1-p^{-j})\geq P_{3}$ is handled by Theorem 11. Hence the decay to zero is quantified precisely. The implied constant is $\approx 0.31$, so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$. The explicit bounds in Section 6 show that standard cryptographic moduli with controlled factorization incur only a small constant-factor overhead; see Remark 1 for effectivity.
3. Limiting distribution (Theorems 6 and 7, Corollary 2, Proposition 4). The quantity $c(p)$ possesses a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the law is moreover purely singular, and Proposition 4 shows that the limiting measure has Hausdorff dimension 0. If $G$ denotes the distribution function, then
$0<G(\tau)<1\qquad(0<\tau<1/2),$
the set of primes with $c(p)\leq\tau$ has relative density $G(\tau)$, and $G(\tau)\to 0$ as $\tau\downarrow 0$. The existence and continuity of $G$ follow from the shifted-prime Erdős–Wintner framework for additive functions evaluated at p1p-1 [2]; for the relevant classical background see Tenenbaum [3] and Deshouillers–Hassani [4]. Pure singularity is classical in the integer case (Erdős [5]) and carries over to shifted primes by the same method, as noted by Tenenbaum [3, p. 263]; we include a self-contained proof for the reader’s convenience. Our new contribution on the distributional side is the full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$, established by combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4). Since the Euler product $P(p)\to 1$, the prim-lwe density $c(p)=\frac{\varphi(p-1)}{p-1}P(p)$ shares the same limiting distribution (Lemma 5). We also derive the moments of $G$ as convergent Euler products (Theorem 8) and establish the analytic continuation of the Mellin transform $\mathbb{E}[X^{s}]$ to the right half-plane (Theorem 10). At the right endpoint, $1-G(\tfrac{1}{2}-\varepsilon)\sim\kappa/\!\log(1/\varepsilon)$ as $\varepsilon\downarrow 0$ for an explicit constant $\kappa$ (Proposition 7), so $G$ is not Hölder of any positive order at $1/2$.
4. Explicit bounds for cryptographic moduli (Theorem 11, Table 1). We give unconditional lower bounds on $c(q)$ for primes $q$ parameterized by $\omega(q-1)$, the number of distinct prime factors of q1q-1, yielding concrete bounds for the NIST-standardized primes and other moduli whose predecessors have controlled factorization structure. The following consequence singles out the cryptographic implication: in the reduction from decision-LWE to decision-prim-lwe over $\mathbb{F}_{q}$, the dimension-uniform expected rejection-sampling overhead $1/c(q)$ admits an explicit upper bound depending only on $\omega(q-1)$.
@@
Let
$P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779.$
Let $q\geq 3$ be a prime modulus with $\omega(q-1)\leq K$. Then
$\frac{1}{c(q)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
The general $\omega(q-1)$-based bound is immediate from Theorem 11; the sharper NIST figures use the exact factorizations recorded in Section 6.
@@
A related distribution problem was studied by Aivazidis and Sofos [6], who considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ as $q$ varies over prime powers and proved corresponding distribution laws. Their setting is different from ours: we work over prime fields and with the primitive-root determinant density arising in prim-lwe. On the distributional side, Deshouillers and Hassani [4, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [4, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even. The analogous support question for the classical distribution of $\varphi(n)/n$ over the integers remains open; see [4].
Organization
Section 2 records the precise definitions and the open question from [1]. Section 3 gives the unconditional proof that $\inf_{p}c(p)=0$. Section 4 proves the upper bound $\min_{p\leq x}c(p)\ll 1/\log\log x$. Section 5 develops the limiting distribution theory, including the Hausdorff-dimension-zero result and the moment Euler product (Section 5.3). Section 6 derives explicit lower bounds for primes of cryptographic interest and, in particular, supplies the matching lower bound $\min_{p\leq x}c(p)\gg 1/\log\log x$. Section 7 discusses the implications for prim-lwe.
Notation
We write $\omega(n)$ for the number of distinct prime factors of an integer $n$, $\pi(x)$ for the number of primes up to $x$, and adopt the convention that $p_{1}=2<p_{2}=3<p_{3}=5<\cdots$ denotes the sequence of all primes in increasing order. For a prime $p$, we write
$P(p):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
for the matrix-count Euler product; thus $c(p)=\frac{\varphi(p-1)}{p-1}\,P(p)$. All limits are taken as $x\to\infty$ unless stated otherwise. We use the standard Vinogradov notation $A\ll B$ and $A\gg B$ synonymously with $A=O(B)$ and $B=O(A)$, respectively. $A\asymp B$ means both $A\ll B$ and $A\gg B$. We write $\operatorname{Li}(x)=\int_{2}^{x}\mathrm{d}t/\!\log t$ for the logarithmic integral, and throughout $\log$ denotes the natural logarithm. We use the notation $A=O^{*}(B)$ to mean $|A|\leq B$, i.e. an implied constant of $1$.
@@ 2 Background and the Open Question
We now record the precise setup from [1] and state the open question. Following [1], define
@@
is a surjective homomorphism, so all its fibres have the same cardinality. Therefore a fraction $\varphi(p-1)/(p-1)$ of invertible matrices have primitive-root determinant. Since
$|\mathrm{GL}_{n}(\mathbb{F}_{p})|=p^{n^{2}}\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
it follows that
@@
recovering [1, Corollary 4]. Consequently, the limit
@@
exists and is strictly positive for each fixed $p$. The quantity $c_{n}(p)$ is the exact acceptance probability in dimension $n$. Since $c_{n}(p)$ is decreasing in $n$, the dimension-uniform constant that enters the prim-lwe reduction is
$\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)=c(p).$
Thus $c(p)$ is not the exact finite-dimensional density but the limiting, dimension-uniform reduction constant.
In the matrix form of decision-LWE, given $\mathbf{A}\in\mathbb{Z}_{q}^{w\times n}$, one must distinguish
$(\mathbf{A},\,\mathbf{A}\mathbf{S}+\mathbf{E})$
from uniform over $\mathbb{Z}_{q}^{w\times n}\times\mathbb{Z}_{q}^{w\times n}$, where $\mathbf{S}\in\mathbb{Z}_{q}^{n\times n}$ is the secret matrix and $\mathbf{E}\in\mathbb{Z}_{q}^{w\times n}$ is the error matrix [7]. Via the identification $\mathbb{Z}/q\mathbb{Z}\cong\mathbb{F}_{q}$ for prime $q$, decision-prim-lwe [1] is the variant in which the secret matrix is additionally required to lie in $M_{n}^{\mathrm{prim}}(\mathbb{F}_{q})$.
In [1, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes
$p=\prod_{i=1}^{k}p_{i}+1,$
then
$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{k}\frac{p_{i}-1}{p_{i}}\longrightarrow 0$
by Mertens’ theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally; see [8] for computational results on primorial primes.
@@ 3 Unconditional Proof that $\inf_{p}c(p)=0$
The argument requires only two classical results.
Theorem 1 (Dirichlet [9], 1837).
For coprime integers $a$ and $q$, there are infinitely many primes $p\equiv a\pmod{q}$.
Theorem 2 (Mertens’ Third Theorem [10], 1874).
@@
$\prod_{\begin{subarray}{c}q\leq x\\ q\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{q}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log x},$
@@
For each $J\geq 1$, set $a_{j}:=p^{-j}$. Then $0<a_{j}\leq 1$ for all $j$, so the finite-product inequality
$\prod_{j=1}^{J}(1-a_{j})\geq 1-\sum_{j=1}^{J}a_{j}$
gives
$\prod_{j=1}^{J}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;1-\sum_{j=1}^{J}\frac{1}{p^{j}}.$
Letting $J\to\infty$ yields
$\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;1-\sum_{j=1}^{\infty}\frac{1}{p^{j}}=1-\frac{1}{p-1}\geq\frac{1}{2}\qquad(p\geq 3).$
Theorem 3 (Unconditional vanishing of the infimal PRIM-LWE density).
@@
Since $\gcd(1,N_{k})=1$, Theorem 1 guarantees infinitely many primes in the arithmetic progression $\{1+j\,N_{k}:j\geq 1\}$. Fix any such prime $p\equiv 1\pmod{N_{k}}$, so that $N_{k}\mid(p-1)$.
Because every prime $q\leq p_{k}$ divides p1p-1, the multiplicative formula for the totient gives
$\frac{\varphi(p-1)}{p-1}\;=\;\prod_{q\mid(p-1)}\Bigl(1-\frac{1}{q}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
where the inequality holds because each factor $\bigl(1-\frac{1}{q}\bigr)$ lies in $(0,1)$, so the full product over all prime divisors $q\mid(p-1)$ is no larger than the partial product over the subset $p_{1},\dots,p_{k}$.
@@ 4 Quantitative Refinements
We first record the form of Linnik’s theorem used in the proof.
Theorem 4 (Least prime in an arithmetic progression; Linnik [11]).
There exist absolute constants $C_{0},L>0$ such that for every modulus $q\geq 1$ and every residue class $a\pmod{q}$ with $\gcd(a,q)=1$, the least prime $p(a,q)\equiv a\pmod{q}$ satisfies
$p(a,q)\leq C_{0}q^{L}.$
Let $N_{k}:=\prod_{i=1}^{k}p_{i}$ denote the $k$-th primorial, $\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x,\;p\text{ prime}\end{subarray}}\log p$ the Chebyshev function, and $m(x):=\max\{k\geq 1:N_{k}\leq x\}$.
Lemma 2 (Prime-number-theorem consequences for primorials; cf. [12]).
@@
The first two assertions are standard consequences of the prime number theorem. Since
$N_{m(x)}\leq x<N_{m(x)+1}=N_{m(x)}\,p_{m(x)+1},$
we have
$\log N_{m(x)}\leq\log x<\log N_{m(x)}+\log p_{m(x)+1}.$
Now
$\log p_{m(x)+1}=o\!\bigl(p_{m(x)}\bigr)=o\!\bigl(\log N_{m(x)}\bigr),$
because $\log N_{m(x)}=\vartheta(p_{m(x)})\sim p_{m(x)}$ and $p_{m(x)+1}\sim p_{m(x)}$. Hence $\log x\sim\log N_{m(x)}$, and therefore $p_{m(x)}\sim\log x$. ∎
@@
By Theorem 4, fix absolute constants $C_{0},L>0$ such that the least prime in any reduced residue class modulo $q$ is at most $C_{0}q^{L}$. For each $k\geq 1$, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$. Given $x$, let $k$ be the largest positive integer such that

### 21 Apr 2026 v4

@@ Abstract
For a prime $p$, let $c_{n}(p)$ denote the fraction of $n\times n$ matrices over $\mathbb{F}_{p}$ whose determinant is a primitive root modulo $p$, and let $c(p)=\lim_{n\to\infty}c_{n}(p)$ be the limiting density. This quantity governs the efficiency of PRIM-LWE, a variant of the learning with errors (LWE) problem in which the secret matrix is required to have primitive-root determinant. The value $1/c(p)$ equals the expected rejection-sampling overhead in the reduction from standard LWE to PRIM-LWE. We prove unconditionally that $\inf_{p\text{ prime}}c(p)=0$, resolving an open question from the PRIM-LWE literature, and establish the rate $\min_{p\leq x}c(p)\asymp 1/\log\log x$ as $x\to\infty$. We show that $c(p)$, viewed as a function on the primes, has a continuous, purely singular limiting distribution with support $[0,1/2]$ and Hausdorff dimension 0. The moments of this limiting distribution are given by convergent Euler products, and the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\operatorname{Re}(s)>0$. Using the number of distinct prime factors $\omega(p-1)$, we derive explicit lower bounds: for every prime $p>2^{30}$ the overhead satisfies $1/c(p)<1.79\log p$, and in general $1/c(p)=O(\log\log p)$. We also describe a certified algorithm for evaluating $c(p)$ and tabulate values for primes of cryptographic interest.
Keywords: Primitive root, determinant density, PRIM-LWE, limiting distribution, Hausdorff dimension
MSC 2020: 11A07, 11N60, 94A60, 15B33
@@ 1 Introduction
Let $p$ be a prime and $M_{n}(\mathbb{F}_{p})$ denote the ring of $n\times n$ matrices over $\mathbb{F}_{p}$. Following [36], a matrix $M\in M_{n}(\mathbb{F}_{p})$ has primitive-root determinant if $\det(M)$ generates $\mathbb{F}_{p}^{*}$. Write $c_{n}(p)$ for the fraction of such matrices, and $c(p):=\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)$ for the limiting density; the precise formulas are recorded in Section 2. Since $c_{n}(p)$ is decreasing in $n$, one has $c(p)\leq c_{n}(p)$ for every $n\geq 1$. Consequently, every lower bound on $c(p)$ is simultaneously a lower bound on every finite-dimensional density $c_{n}(p)$.
The quantity $c(p)$ arises as a reduction constant in the PRIM-LWE variant [36] of the popular learning with errors (LWE) problem [33], introduced to construct verifiable secret sharing schemes with “free” verification. The reduction proceeds by rejection-sampling and succeeds with probability $c_{n}(p)$ in dimension $n$, so the dimension-uniform overhead is $1/c(p)$. This overhead represents a tightness loss in the security reduction: the hardness guarantee for PRIM-LWE degrades by a factor of $c(p)$ relative to standard LWE, making tight bounds on $c(p)$ essential for concrete security estimates. Sehrawat, Yeo, and Desmedt [36, Remark 3(ii)] asked whether $\inf_{p}c(p)=0$ can be proved unconditionally.
@@
2. Sharp order of the minimum (Proposition 1, Corollary 7). We establish
@@
See Remark 4 for effectivity.
3. Limiting distribution (Theorems 4 and 5, Corollary 2, Proposition 4). The quantity $c(p)$ possesses a continuous, purely singular limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the limiting measure has Hausdorff dimension 0. The full-support conclusion is new; it combines the left non-differentiability result of Deshouillers–Hassani [9, Theorem 1.1] with the density of even totient ratios (Lemma 4). The moments of $G$ are given by convergent Euler products (Theorem 6), the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\{\operatorname{Re}(s)>0\}$ (Theorem 8), and $G$ is not Hölder of any positive order at $1/2$ (Proposition 7).
Since the limiting law is purely singular and supported on a set of Hausdorff dimension 0, it cannot certify $c(p)$ at any individual prime; this motivates the pointwise bounds of item 4.
4. Explicit bounds via distinct prime factors (Theorem 9). We give unconditional lower bounds on $c(p)$ for primes $p$ parameterized by $\omega(p-1)$, the number of distinct prime factors of p1p-1, yielding computable bounds for any prime $p$ such that the factorization of p1p-1 is known. For instance, for the NIST-standardized primes $p=3329$ and $p=8380417$, the exact factorizations of p1p-1 give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$; see Section 7. In particular, the dimension-uniform expected rejection-sampling overhead $1/c(p)$ in the LWE-to-PRIM-LWE reduction [36, Theorem 10] admits an explicit upper bound depending only on $\omega(p-1)$.
From the computational viewpoint, our main contributions are: explicit lower bounds on $c(p)$ valid for all primes $p>2^{30}$ (Corollary 9), a certified rational-arithmetic algorithm for evaluating $c(p)$ to arbitrary precision given the factorization of p1p-1 (Proposition 8), and tables of $c(p)$ for all primes appearing in current NIST post-quantum cryptographic standards and representative lattice-based implementations (Tables 2 and 3). The explicit bound takes the following compact form:
@@
Let $P_{3}:=\prod_{j=1}^{\infty}(1-1/3^{j})\approx 0.5601260779$. If $p\geq 3$ is a prime modulus with $\omega(p-1)\leq K$, then
$\frac{1}{c(p)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
The general $\omega(p-1)$-based bound is immediate from Theorem 9. Table 1 lists representative values of $B(K)$, and Tables 2 and 3 collect these bounds for representative cryptographic moduli.
@@
A related distribution problem was studied by Aivazidis and Sofos [1], who considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ as $q$ varies over prime powers and proved corresponding distribution laws. Their setting is different from ours: we work over prime fields and with the primitive-root determinant density arising in PRIM-LWE. On the distributional side, Deshouillers and Hassani [9, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [9, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even. The analogous support question for the classical distribution of $\varphi(n)/n$ over the integers remains open; see [9].
@@ 2 Background and the Open Question
Throughout, $\log$ denotes the natural logarithm. We write $\varphi$ for Euler’s totient function, $\omega(n)$ for the number of distinct prime factors of $n$, and $p_{k}$ for the $k$-th prime.
Following [36], define
@@
is a surjective group homomorphism, so a fraction $\varphi(p-1)/(p-1)$ of invertible matrices have primitive-root determinant. Since
$|\mathrm{GL}_{n}(\mathbb{F}_{p})|=p^{n^{2}}\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
(see, e.g., Mullen and Panario [30, Theorem 13.3.3]), it follows that
@@
cf. [36, Corollary 4]. Consequently, the limit
@@
exists and is strictly positive for each fixed $p$. Since $c_{n}(p)$ is decreasing in $n$, $\inf_{n\geq 1}c_{n}(p)=c(p)$.
Writing $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$ for the prime factorization of p1p-1, one has
$c(p)\;=\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{r_{i}}\Bigr)\;\cdot\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr).$ (1)
Since $P(p):=\prod_{j=1}^{\infty}(1-p^{-j})\to 1$ as $p\to\infty$, the behavior of $c(p)$ over the primes is governed by $\varphi(p-1)/(p-1)$, and the distributional theory in Section 5 reduces to the classical shifted-prime distribution of this ratio.
In [36, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes $p=\prod_{i=1}^{k}p_{i}+1$, then $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(p_{i}-1)/p_{i}\to 0$ by Mertens’ third theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally; see [7] for computational results on primorial primes.
@@ 3 Unconditional Proof that $\inf_{p}c(p)=0$
The argument requires only two classical results: Dirichlet’s theorem on primes in arithmetic progressions [10] and the following.
Theorem 1 (Mertens’ Third Theorem [28]; see [29, Theorem 2.7(e)], 1874).
@@
$\prod_{\begin{subarray}{c}\ell\leq x\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{\ell}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log x},$
@@
By the finite-product inequality $\prod(1-a_{j})\geq 1-\sum a_{j}$ (immediate by induction when $0\leq a_{j}\leq 1$) and the geometric series $\sum_{j\geq 1}p^{-j}=1/(p-1)\leq 1/2$. ∎
Theorem 2 (Unconditional vanishing of $\inf c(p)$).
The following identities hold:
@@
Since $\gcd(1,N_{k})=1$, Dirichlet’s theorem guarantees infinitely many primes in the arithmetic progression $\{1+j\,N_{k}:j\geq 1\}$. Fix any such prime $p\equiv 1\pmod{N_{k}}$, so that $N_{k}\mid(p-1)$.
Because every prime $\ell\leq p_{k}$ divides p1p-1, the multiplicative formula for the totient gives
$\frac{\varphi(p-1)}{p-1}\;=\;\prod_{\ell\mid(p-1)}\Bigl(1-\frac{1}{\ell}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
@@ 4 Quantitative Refinements
By Linnik’s theorem [26], there exist absolute constants $C_{0},L>0$ such that the least prime in any reduced residue class modulo $d$ is at most $C_{0}d^{L}$.
Let $N_{k}$ denote the $k$-th primorial, $\vartheta$ the Chebyshev function, and $m(x)$ the primorial-counting function:
$N_{k}:=\prod_{i=1}^{k}p_{i},\qquad\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x\\ p\text{ prime}\end{subarray}}\log p,\qquad m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
Lemma 2 (Prime-number-theorem consequences for primorials; cf. [29, Corollary 2.5]).
@@
The first two assertions follow from $\vartheta(x)\sim x$. For the last two, $N_{m(x)}\leq x<N_{m(x)}p_{m(x)+1}$ gives $\log x\sim\log N_{m(x)}\sim p_{m(x)}$, since $\log p_{m(x)+1}=o(p_{m(x)})$. ∎
@@
Fix absolute constants $C_{0},L>0$ from Linnik’s theorem such that the least prime in any reduced residue class modulo $d$ is at most $C_{0}d^{L}$. For each $k\geq 1$, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$. Given $x$, let $k$ be the largest positive integer such that
@@
The matching lower bound $\min_{p\leq x,\,p\text{ prime}}c(p)\gg 1/\!\log\log x$, which completes the proof that the order is sharp, is established in Corollary 7 below; its proof requires only the primorial counting argument and Lemma 1.